OpenStudy (gabbyalicorn):

Four friends played mini golf. Their scores were –3, –2, –1, and 2. What is the mean absolute deviation of their scores? step 1 = 2 step 2 = -3- 2 = 1 -2-2= 0 -1-2= 3 2-3=1 i feel like wrong though, like the negatives are tricking me or something :/ @black_ninja123

OpenStudy (black_ninja123):

Step 1 is wrong

OpenStudy (gabbyalicorn):

oh, :/ darn let me try again.

OpenStudy (gabbyalicorn):

wait its kind of harder when its negatives >:P grrr

OpenStudy (black_ninja123):

Use a calc

OpenStudy (black_ninja123):

If you are allow to

OpenStudy (gabbyalicorn):

calculous!!?? o.o i dont no how to do that

OpenStudy (black_ninja123):

Calculator

OpenStudy (gabbyalicorn):

im in middle school not high school

OpenStudy (black_ninja123):

lol calc means calculator

OpenStudy (gabbyalicorn):

oh, sorry

OpenStudy (black_ninja123):

Its fine

OpenStudy (here_to_help15):

Need help?

OpenStudy (gabbyalicorn):

yeppers

OpenStudy (gabbyalicorn):

i got 4

OpenStudy (gabbyalicorn):

wait no! not done yet!

OpenStudy (gabbyalicorn):

i got -1

OpenStudy (black_ninja123):

Correct

OpenStudy (gabbyalicorn):

:D yay!

OpenStudy (black_ninja123):

Now find the difference

OpenStudy (gabbyalicorn):

now the hard part

OpenStudy (gabbyalicorn):

-3 - -1 = -2 ?

OpenStudy (black_ninja123):

Yes

OpenStudy (gabbyalicorn):

omg yasss

OpenStudy (black_ninja123):

But remember absolute deviation

OpenStudy (gabbyalicorn):

-2 - (-2) = 0

OpenStudy (black_ninja123):

Yes

OpenStudy (gabbyalicorn):

hmm kay -1 - (-2)= 1 ?

OpenStudy (black_ninja123):

Correct

OpenStudy (gabbyalicorn):

what!? srsly!? omg i thought i just totally would fail that 1 XD

OpenStudy (gabbyalicorn):

2 - (-2) = 0 ?

OpenStudy (black_ninja123):

No its 4

OpenStudy (black_ninja123):

remember when there is a negative and a positive you have to add

OpenStudy (gabbyalicorn):

kk √

OpenStudy (gabbyalicorn):

so then 4 0 1 -2 -1 div by 5 = 0.4

OpenStudy (black_ninja123):

remove the negative sign from -2 and -1

OpenStudy (gabbyalicorn):

7.2

OpenStudy (black_ninja123):

No

OpenStudy (gabbyalicorn):

D'':

OpenStudy (black_ninja123):

can you tell me what answer you got for step 2

OpenStudy (gabbyalicorn):

o-o all those numbers 4 0 1 2 1

OpenStudy (black_ninja123):

How did you get the 4

OpenStudy (gabbyalicorn):

oh good question..... i guess well take tht out

OpenStudy (black_ninja123):

yea remove the 4 and put a 3

OpenStudy (gabbyalicorn):

1.4

OpenStudy (gabbyalicorn):

? correctamoondo?

OpenStudy (black_ninja123):

One sec checking

OpenStudy (gabbyalicorn):

kk take your time ^u^

OpenStudy (black_ninja123):

its 1.5

OpenStudy (gabbyalicorn):

darnit, so close

OpenStudy (gabbyalicorn):

0.1 off

OpenStudy (black_ninja123):

I think its 1.4 too one sec let me double check

OpenStudy (gabbyalicorn):

kks

OpenStudy (black_ninja123):

@Here_to_Help15

OpenStudy (black_ninja123):

Do you have any answer choices?

OpenStudy (gabbyalicorn):

nope, just fill ins ;-;

OpenStudy (gabbyalicorn):

i hate fill ins...

OpenStudy (black_ninja123):

Yep

OpenStudy (gabbyalicorn):

@Nnesha please help...

OpenStudy (black_ninja123):

How did you get 5 numbers 4 0 1 2 1

OpenStudy (black_ninja123):

There should only be 4

OpenStudy (gabbyalicorn):

well can we count the zero out? probs not :/

OpenStudy (gabbyalicorn):

then u said the 4 is a 3

OpenStudy (black_ninja123):

-3 - (-1)=-2 -2-(-1)=1 -1-(-1)=0 2-(-1)=3

OpenStudy (black_ninja123):

ignore the negative sign

OpenStudy (gabbyalicorn):

kk

OpenStudy (black_ninja123):

2+1+0+3=6 6/4=1.5

OpenStudy (black_ninja123):

1.5 is the answer

OpenStudy (gabbyalicorn):

Okay, Thank you sooo much!

OpenStudy (black_ninja123):

Your welcome Do you need more help?

OpenStudy (gabbyalicorn):

oh im good thank you ^u^